# Generative Point Cloud Registration

### v1

@@ 3.4 Few-Shot Consistency Fine-tuning
Table 1: Comparison of the methods on rotation, translation, and Chamfer distance on ScanNet (dai2017scannet) benchmark dataset.
@@
FPFH (rusu2009fast) 41.4 56.7 73.3 39.2 7.1 17.5 35.1 50.9 79.5 23.5 32.3 48.0 53.0 159.6 6.5
Lepard (li2022lepard) 63.3 75.5 84.1 24.9 3.3 31.3 56.4 72.3 48.4 8.1 51.6 69.1 73.0 89.2 0.9
RegTR (yew2022regtr) 72.5 83.8 94.1 10.2 2.3 44.3 65.6 80.0 27.7 5.8 61.0 76.6 80.9 54.0 0.5
RoITr (yu2023rotation) 70.0 77.5 83.7 24.1 2.3 40.3 62.3 75.1 45.6 6.5 58.8 72.4 75.4 94.1 0.6
FCGF (choy2019fully) 78.9 84.2 87.5 19.4 1.5 55.3 70.7 79.7 37.8 4.3 67.3 78.2 80.3 100.7 0.4
@@
Table 2: Comparison of the methods on rotation, translation, and Chamfer distance on 3DMatch (zeng20173dmatch) benchmark dataset.
@@
FPFH (rusu2009fast) 69.1 82.9 91.2 15.0 3.1 25.8 53.9 75.1 37.4 9.1 52.5 74.2 79.2 57.6 0.9
Lepard (li2022lepard) 84.3 91.0 94.1 11.1 2.1 43.1 75.2 88.9 21.8 5.8 72.1 88.3 90.5 45.3 0.4
RegTR (yew2022regtr) 86.2 92.1 97.2 5.7 1.6 55.0 77.6 88.9 18.8 4.6 75.4 88.2 91.3 40.0 0.3
RoITr (yu2023rotation) 86.3 91.1 93.8 11.1 1.6 51.2 77.4 89.1 20.5 4.9 75.2 88.5 90.6 50.1 0.4
FCGF (choy2019fully) 90.4 93.7 94.8 9.4 1.4 53.4 79.3 91.0 19.2 4.7 76.7 90.8 92.4 40.3 0.4
@@ 3.5 Geometric-Color Fused Point Descriptor
In this section, we focus on how to enhance the geometric representations of point clouds with the free-lunch color information from generated source and target images, $\tilde{\mathbf{I}}_{\mathcal{P}}$ and $\tilde{\mathbf{I}}_{\mathcal{Q}}$. We provide two geometric-color fusion schemes:
Zero-Shot Geometric-Color Feature Fusion. Inspired by the powerful RGB representations of large vision models, we utilize them to directly extract zero-shot semantic features from the generated images. Specifically, we employ two widely-used vision foundation models: DINOv2 (oquab2023dinov2) and Stable Diffusion (rombach2022high) for image encoding, achieving corresponding feature maps. These feature maps are then projected into the point cloud space using the camera intrinsic matrix, yielding pointwise color descriptors: $\{\mathbf{f}^{rgb}_{\mathbf{p}_{i}}\}$ and $\{\mathbf{f}^{rgb}_{\mathbf{q}_{i}}\}$ for both source and target point clouds. Finally, we apply a simple weighted concatenation to combine the RGB descriptors with the geometric descriptors, resulting in fused descriptors as follows:
@@
where $\cdot$$\cdot$] denotes the feature concatenation operator; $\omega$ is the fusion weight and $\mathbf{f}_{\mathbf{p}_{i}}^{geo}$ represents the geometric descriptors; $\operatorname{PCA}(\cdot)$ denotes the principal component analysis function to compress the feature dimension of the color descriptor to fit that of the geometric descriptor. The same fusion scheme is also applied to the target point clouds. Notably, this zero-shot geometric-color fusion approach is general and can be applied to a variety of geometric descriptors, whether traditional or deep descriptors.
XYZ-RGB Fusion. This fusion scheme directly projects the generated source and target RGB images into the point cloud space. The resulting RGB values are then concatenated with the point coordinates of the point clouds, forming 6D color source and target point clouds, as shown in Fig. 5 (left). These color point clouds are subsequently used as inputs to the color point cloud registration method, like ColorPCR (mu2024colorpcr), for 3D registration.
Figure 5: Left: The visualization of the generated RGB image pairs and the formed color source and target point clouds; Right: In low-overlap cases, the original Predator struggles with registration. By contrast, the Generative Predator, enhanced with generated color information, successfully align them well.
@@ 4.1 Experimental Setting
Implementation Details. During the few-shot fine-tuning stage, we randomly select 3,000 sample pairs from the ScanNet training set (dai2017scannet) for model fine-tuning. Following the default fine-tuning configuration of ControlNet (Zhang_2023_ICCV), we adopt the AdamW optimizer (loshchilov2017decoupled) with a learning rate of 1e-5 and set the training epoch to 10. The code for this project is implemented in PyTorch, and all experiments are conducted on a server equipped with an Intel i5 2.2 GHz CPU and a TITAN RTX GPU. In our experiments, we integrate our zero-shot geometric-color fusion (Sec. 3.5) with three prevalent deep geometric descriptors: FCGF (choy2019fully), Predator (huang2021predator), and GeoTransformer (qin2022geometric), resulting in corresponding color-enhanced variants: Generative FCGF, Generative Predator, and Generative GeoTrans for method evaluation. Additionally, to validate our XYZ-RGB fusion scheme, we replace the real color point clouds (with actual RGB values) used by ColorPCR (mu2024colorpcr) with our generated color point clouds (with synthesized RGB values), forming Generative ColorPCR for 3D matching.
Evaluation Metric. Following (el2021unsupervisedr; yuan2023pointmbf), we use rotation error, translation error, and Chamfer error, including the accuracy across varying thresholds and mean/median errors, for performance evaluation.
@@ 4.2 Comparison with Existing Methods
Evaluation on ScanNet. We first perform model evaluation on a widely-used, large-scale indoor benchmark dataset, ScanNet (dai2017scannet). We follow the official data split to divide this dataset into the training, validation, and testing subsets, and construct view pairs by sampling image pairs that are 50 frames apart. Compared to the 20-frame separation used in (el2021unsupervisedr; yuan2023pointmbf), our approach with a 50-frame separation further reduces the overlap ratio (i.e., lower overlap), thereby increasing the registration difficulty. We compare our method against with one representative traditional descriptor: FPFH (rusu2009fast), one scene-level end-to-end registration network: RegTR (yew2022regtr), and five deep descriptors: FCGF (choy2019fully), Predator (huang2021predator), GeoTrans (qin2022geometric), Lepard (li2022lepard), and RoITr (yu2023rotation). We adopt RANSAC-50k as the pose estimator for FPFH, Lepard and RoITr, and select SC2PCR (chen2022sc2), RANSAC-50k and LGR for (Generative) FCGF, (Generative) Predator, and (Generative) GeoTrans (qin2022geometric), to validate the robustness of our generative 3D registration paradigm across different pose estimators. Table 1 demonstrates that enhanced by the free-lunch color information generated by our Match-ControlNet, all generative versions of FCGF, Predator, and GeoTrans achieve significant performance improvements, such as $6.9\%\uparrow$ of Generative FCGF on 45@Rotation metric. These confirm the generality and effectiveness of our proposed generative point cloud registration paradigm. Additionally, we find that compared to the DINOv2 image encoding, Stable Diffusion can capture more discriminative representations and achieve higher precisions.
Evaluation on 3DMatch. We next evaluate our method on 3DMatch (zeng20173dmatch), another widely-used benchmark dataset for 3D registration. We follow (el2021unsupervisedr; yuan2023pointmbf) as in ScanNet to produce the pairwise samples. Also, we increase the view separation from 20 to 40, resulting in point cloud pairs with lower overlap to increase the registration challenge. Table. 2 demonstrates that by incorporating FCGF, Predator, and GeoTrans into our generative point cloud registration framework, their generative variants also consistently achieve the performance gain, validating the effectiveness of our proposed paradigm.
Table 3: Ablation studies on 3DMatch (zeng20173dmatch) dataset. (*) denotes the default configuration.
@@ 4.3 Ablation Studies and Analysis
Effectiveness of Match-ControlNet. We first evaluate the performance contribution of our developed Match-ControlNet: (i) The top block of Table 3 demonstrates that, compared to using generated image pairs with only 2D-3D geometric consistency (geo), incorporating both 2D-3D geometric consistency and cross-view texture consistency (geo+tex) through our Match-ControlNet results in higher registration accuracy. This improvement is due to the additional benefit of consistent textures and colors, which further facilitate accurate correspondence identification. Additionally, we observe that the generated images with only 2D-3D geometric consistency can also bring performance gain in some criteria. We attribute it to that DINOv2 and Stable Diffusion can extract powerful semantic representations, mitigating the feature inconsistency of correspondences caused by the texture difference and thereby aiding correspondence identification. Furthermore, we visualize the generated image pair for given source and target point clouds in Fig. 5 (left). It shows that our Match-ControlNet is capable of producing high-quality image pairs with consistent 2D-3D geometry and cross-view texture.
@@
Zero-Shot vs Finetuning. We further investigate the performance of Match-ControlNet in both zero-shot and finetuned settings. As shown in the second block of Table 3, both approaches yield substantial improvements over FCGF. Moreover, because the finetuned Match-ControlNet benefits from task-specific training, it consistently achieves higher registration accuracy than the zero-shot version. Notably, even few-shot finetuning with as few as 1K samples yields clear performance gains. Increasing the number of finetuning samples (e.g., to 3K or 5K) provides additional improvements; however, models trained on 3K or 5K samples show comparable registration accuracy in practice. Hence, we adopt 3K samples as our default finetuning configuration.
@@
Zero-Shot Geometric-Color Feature Fusion. We next conduct ablation studies on the zero-shot geometric-color feature fusion described in Eq. 6. As shown in the fourth block of Table 3, Generative GeoTrans exhibits varying registration performance under different color feature dimensions, $d_{rgb}\in\{16,32,64,128\}$. We observe that a very small color feature dimension (e.g., $d_{rgb}=16$) degrades performance due to limited semantic representational capacity, while excessively large dimensions do not yield significant performance gains. Therefore, to balance inference efficiency with registration precision, we set $d_{rgb}=64$ as our default setting. Additionally, in the fifth block of Table 3, we investigate performance variations with different fusion weights $\omega\in\{0.0,0.25,0.50,0.75,1.0\}$, where a larger $\omega$ places more emphasis on the geometric descriptors (see Eq. 6). Our results indicate that both overly high $\omega$ (which overemphasizes geometry) and overly low $\omega$ (which overemphasizes color) lead to degraded registration accuracy. By contrast, a balanced weight (e.g., $\omega=0.50$) achieves higher performance. As a result, we adopt $\omega=0.50$ as our default hyperparameter configuration.
Figure 6: Our Match-ControlNet effectively mitigates calibration errors and lighting challenges commonly encountered in real-world RGB-D data, thereby improving the matching precision of color point cloud registration methods.
@@
XYZ-RGB Fusion. We finally evaluate the effectiveness of the XYZ-RGB fusion (see Sec. 3.5). The third block in Table 3 demonstrates that, on the 3DMatch dataset, Generative ColorPCR with the synthesized color even outperforms the original ColorPCR with the real color. This advantage is attributed not only to the high-quality pairwise image generation provided by our Match-ControlNet, but also to several key benefits of our generated XYZ-RGB data over real XYZ-RGB data: (i) Mitigating calibration errors: As shown in Fig. 6 (left), some real RGB-D data would suffer from calibration errors, which may lead to misalignment in the colored point clouds. By contrast, our framework, benefiting from the powerful 2D-3D consistency generation ability, effectively reduces such calibration errors, producing higher-quality colored point clouds and enabling more accurate matching. (ii) Mitigating lighting challenges: Fig. 6 (right) shows that RGB data from real-world conditions can degrade under poor lighting, negatively impacting RGB-D matching performance. Our generative point cloud registration framework, however, can generate images with consistent lighting conditions, independent of real-world lighting issues, thus enhancing the overall lighting robustness.

### v2

@@ 3.5 Theoretical Analysis of Coupled Denoising Model
TABLE II: Comparison of the methods on rotation, translation, and Chamfer distance on ScanNet [11] benchmark dataset.
@@
FPFH [34] 41.4 56.7 73.3 39.2 7.1 17.5 35.1 50.9 79.5 23.5 32.3 48.0 53.0 159.6 6.5
Lepard [25] 63.3 75.5 84.1 24.9 3.3 31.3 56.4 72.3 48.4 8.1 51.6 69.1 73.0 89.2 0.9
RegTR [49] 72.5 83.8 94.1 10.2 2.3 44.3 65.6 80.0 27.7 5.8 61.0 76.6 80.9 54.0 0.5
RoITr [51] 70.0 77.5 83.7 24.1 2.3 40.3 62.3 75.1 45.6 6.5 58.8 72.4 75.4 94.1 0.6
FCGF [10] 78.9 84.2 87.5 19.4 1.5 55.3 70.7 79.7 37.8 4.3 67.3 78.2 80.3 100.7 0.4
@@
TABLE III: Comparison of the methods on rotation, translation, and Chamfer distance on 3DMatch [54] benchmark dataset.
@@
FPFH [34] 69.1 82.9 91.2 15.0 3.1 25.8 53.9 75.1 37.4 9.1 52.5 74.2 79.2 57.6 0.9
Lepard [25] 84.3 91.0 94.1 11.1 2.1 43.1 75.2 88.9 21.8 5.8 72.1 88.3 90.5 45.3 0.4
RegTR [49] 86.2 92.1 97.2 5.7 1.6 55.0 77.6 88.9 18.8 4.6 75.4 88.2 91.3 40.0 0.3
RoITr [51] 86.3 91.1 93.8 11.1 1.6 51.2 77.4 89.1 20.5 4.9 75.2 88.5 90.6 50.1 0.4
FCGF [10] 90.4 93.7 94.8 9.4 1.4 53.4 79.3 91.0 19.2 4.7 76.7 90.8 92.4 40.3 0.4
@@
Figure 5: The visualization of the generated perspective-view/panoramic RGB image pairs and the formed color source and target point clouds using the DepthMatch-ControlNet (left) and LiDARMatch-ControlNet (right).
@@ 3.6 Geometric-Color Fused Point Descriptor
In this section, we focus on how to enhance the geometric representations of point clouds with the free-lunch color information from generated perspective/panoramic source and target images, denoted as $\tilde{\mathbf{I}}_{\mathcal{P}}$ and $\tilde{\mathbf{I}}_{\mathcal{Q}}$. We provide two geometric-color fusion schemes:
Zero-Shot Geometric-Color Feature Fusion. Inspired by the powerful RGB representations of large vision models, we utilize them to directly extract zero-shot semantic features from the generated images. Specifically, we employ two widely-used vision foundation models: DINOv2 [31] and Stable Diffusion [33] for image encoding, achieving corresponding feature maps. These feature maps are then projected into the point cloud space using the camera intrinsic matrix, yielding pointwise color descriptors: $\{\mathbf{f}^{rgb}_{\mathbf{p}_{i}}\}$ and $\{\mathbf{f}^{rgb}_{\mathbf{q}_{i}}\}$ for both source and target point clouds. Finally, we apply a simple weighted concatenation to combine the RGB descriptors with the geometric descriptors, resulting in fused descriptors:
@@
where $\cdot$$\cdot$] denotes the feature concatenation operator; $\omega\in[0,1]$ is the fusion weight and $\mathbf{f}_{\mathbf{p}_{i}}^{geo}$ represents the geometric descriptors; $\operatorname{PCA}(\cdot)$ denotes the principal component analysis function to compress the feature dimension of the color descriptor to fit that of the geometric descriptor. The same fusion scheme is also applied to the target point clouds. Notably, this zero-shot geometric-color fusion approach is general and can be applied to a variety of geometric descriptors, whether traditional or deep descriptors.
XYZ-RGB Fusion. This fusion scheme directly projects the generated source and target RGB images into the point cloud space. The resulting RGB values are then concatenated with the point coordinates of the point clouds, forming 6D color source and target point clouds, as shown in Fig. 5. These color point clouds are subsequently used as inputs to the color point cloud registration method, like ColorPCR [30], for 3D registration.
Figure 6: Qualitative comparions between Predator and Generative Predator on 3DMatch [54]. The original Predator struggles with accurate registration. By contrast, the Generative Predator, enhanced with our generated color information, achieves significantly better alignment.
TABLE IV: Comparison of the methods on FMR, IR, and $R$ on Dur360BEV [52] dataset with point cloud pairs separated by at least 5 or 10 meters.
$\geq$ 5m Apart $\geq$ 10m Apart
Methods FMR (%) $\uparrow$ IR (%) $\uparrow$ RR (%) $\uparrow$ FMR (%) $\uparrow$ IR (%) $\uparrow$ RR (%) $\uparrow$
PARE-Net [47] 93.3 24.3 69.6 76.5 13.5 56.7
CoFiNet [50] 100.0 34.2 99.4 95.6 20.5 90.5
FPFH [34] 75.1 6.4 11.9 14.6 3.8 0.7
Generative FPFHDINOv2 100.0 18.8 86.7 84.9 8.5 30.6
Generative FPFHSD 100.0 21.8 88.6 86.7 9.0 31.0
Improvement $\uparrow$ 24.9 15.4 76.7 72.1 5.2 30.3
FCGF [10] 100.0 41.3 96.3 100.0 22.8 85.3
[…]
Generative FCGFSD 100.0 44.2 98.7 99.9 23.0 91.2
Improvement $\uparrow$ 0.0 2.9 2.4 0.1 1.2 5.9
Predator [21] 99.9 23.0 95.8 97.3 12.6 57.9
[…]
Generative PredatorSD 100.0 33.3 98.4 98.7 16.1 77.7
Improvement $\uparrow$ 0.1 10.3 2.6 1.4 3.5 19.8
GeoTrans [32] 100.0 62.9 99.8 98.3 40.1 94.9
[…]
Generative GeoTransSD 100.0 63.8 99.9 99.8 41.1 95.5
Improvement $\uparrow$ 0.0 0.7 0.2 1.6 1.0 0.9
@@ 4.1 Experimental Setting
Implementation Details. During the few-shot fine-tuning stage, we randomly select 3,000 sample pairs from the ScanNet training set [11] for model fine-tuning. Following the default fine-tuning configuration of ControlNet [55], we adopt the AdamW optimizer [27] with a learning rate of 1e-5 and set the training epoch to 10. The code for this project is implemented in PyTorch, and all experiments are conducted on a server equipped with an Intel i5 2.2 GHz CPU and a TITAN RTX GPU.
Figure 7: Correspondence visualization of Predator and Generative Predator on 3DMatch dataset [54] (top) and Dur360DEV dataset [52] (bottom). Benefitting from the rich RGB information from generated perspective-view/panoramic images, our Generative-Predator descriptors presents higher inlier ratios than the geometry-only Predator descriptors.
Evaluation Metric. Following [13, 53], we use rotation error, translation error, and Chamfer error, including the accuracy across varying thresholds and mean/median errors, for performance evaluation on depth camera-based 3D registration. Instead, for LiDAR-based 3D registration, we follow the evaluation protocol proposed in [21, 32], adopting three metrics: Inlier Ratio (IR), Feature Matching Recall (FMR), and Registration Recall (RR) for evaluation.
@@ 4.2 Comparison with Existing Methods
Evaluation on ScanNet. We first test our generative point cloud registration paradigm on the depth camera-based 3D registration. We perform model evaluation on a widely-used, large-scale indoor benchmark dataset, ScanNet [11], where the point clouds are derived from depth maps captured by depth sensors. We follow the official data split to divide this dataset into the training, validation, and testing subsets, and construct view pairs by sampling image pairs that are 50 frames apart. Compared to the 20-frame separation used in [13, 53], our approach with a 50-frame separation further reduces the overlap ratio (i.e., lower overlap), thereby increasing the registration difficulty.
We compare our method against with one representative traditional descriptor: FPFH [34], one scene-level end-to-end registration network: RegTR [49], and five deep descriptors: FCGF [10], Predator [21], GeoTrans [32], Lepard [25], and RoITr [51]. Here, we integrate FCGF, Predator, and GeoTrans into our generative point cloud registration framework, resulting in corresponding color-enhanced variants: Generative FCGF, Generative Predator, and Generative GeoTrans for method evaluation. We adopt RANSAC-50k as the pose estimator for FPFH, Lepard and RoITr, and select SC2PCR [7], RANSAC-50k and LGR for (Generative) FCGF, (Generative) Predator, and (Generative) GeoTrans [32], respectively, to validate the robustness of our generative 3D registration paradigm across different pose estimators. Table II demonstrates that enhanced by the free-lunch color information generated by our DepthMatch-ControlNet, all generative versions of FCGF, Predator, and GeoTrans achieve significant performance improvements, such as $6.9\%\uparrow$ of Generative FCGF on 45@Rotation metric. These confirm the generality and effectiveness of our framework. Additionally, we find that compared to the DINOv2 image encoding, Stable-Diffusion (SD) encoding can capture more discriminative representations and achieve higher precisions.
Evaluation on 3DMatch. We next evaluate our method on 3DMatch [54], another widely-used benchmark dataset for single-view 3D registration. We follow [13, 53] as in ScanNet to produce the pairwise samples. Also, we increase the view separation from 20 to 40, resulting in point cloud pairs with lower overlap to increase the registration challenge. Table III shows that integrating FCGF, Predator, and GeoTrans into our generative point cloud registration framework consistently improves their performance, further validating the effectiveness of our proposed paradigm. Additional qualitative comparisons and correspondence visualizations are presented in Fig. 6 and Fig. 7, respectively.
TABLE V: Ablation studies on 3DMatch [54] dataset. (*) denotes the default configuration.
@@
Evaluation on Dur360BEV. We evaluate the effectiveness of our generative point cloud registration approach on LiDAR-based 3D registration using outdoor LiDAR-scanned driving scenarios from the Dur360BEV dataset [52]. Following the official data protocol, we use 14,767 frames with timestamps 0–11,509 and 13,150–16,406 as the training set, and the remaining 1,640 frames with timestamps 11,510–13,149 as the testing set. Consistent with [10], we select only point cloud pairs that are at least 5 or 10 meters apart in spatial distance for training and testing, and apply ICP to refine the ground-truth rigid transformations. We compare our method with a traditional descriptor, FPFH [34], and five deep learning-based descriptors: PARE-Net [47], CoFiNet [50], FCGF [10], Predator [21], and GeoTrans [32]. Note that methods such as Lepard, RegTR, and RoITr are not included in the comparison, as their pretrained models are not available for LiDAR settings. We integrate FPFH, FCGF, Predator, and GeoTrans into our generative point cloud registration framework, resulting in corresponding color-enhanced variants: Generative FPFH, Generative FCGF, Generative Predator, and Generative GeoTrans for method evaluation. As shown in Table IV, the proposed generative variants tend to consistently achieve performance gains over their respective baselines, benefitting from the high-quality panoramic image generation provided by our LiDARMatch-ControlNet. Notably, Generative FPFH achieves a Registration Recall (RR) gain of 76.7%, while Generative Predator reaches an Inlier Ratio (IR) gain of 10.3%. These impressive accuracy gains highlight the strong effectiveness of our generative paradigm. More qualitative comparisons are also provided in Fig. 7 and Fig. 8.
@@ 4.3 Ablation Studies and Analysis
Effectiveness of Match-ControlNet. We first evaluate the performance contribution of our developed Match-ControlNet: (i) The top block of Table V demonstrates that, compared to using generated image pairs with only 2D-3D geometric consistency (geo), incorporating both 2D-3D geometric consistency and cross-view texture consistency (geo+tex) through our Match-ControlNet results in higher registration accuracy. This improvement is due to the additional benefit of consistent textures and colors, which further facilitate accurate correspondence identification. Additionally, we observe that the generated images with only 2D-3D geometric consistency can also bring performance gain in some criteria. We attribute it to that DINOv2 and Stable Diffusion can extract powerful semantic representations, mitigating the feature inconsistency of correspondences caused by the texture difference and thereby aiding correspondence identification. Furthermore, we visualize the generated image pair for given source and target point clouds in Fig. 5. It shows that our Match-ControlNet is capable of producing high-quality image pairs with consistent 2D-3D geometry and cross-view texture.
@@
Zero-Shot vs Finetuning. We further investigate the performance of DepthMatch-ControlNet in both zero-shot and finetuned settings. As shown in the second block of Table V, both approaches yield substantial improvements over FCGF. Moreover, because the finetuned DepthMatch-ControlNet benefits from task-specific training, it consistently achieves higher registration accuracy than the zero-shot version. Notably, even few-shot finetuning with as few as 1K samples yields clear performance gains. Increasing the number of finetuning samples (e.g., to 3K or 5K) provides additional improvements; however, models trained on 3K or 5K samples show comparable registration accuracy in practice. Hence, we adopt 3K samples as our default finetuning configuration.
@@
Zero-Shot Geometric-Color Feature Fusion. We next conduct ablation studies on the zero-shot geometric-color feature fusion described in Eq. 12. As shown in the fourth block of Table V, Generative GeoTrans exhibits varying registration performance under different color feature dimensions, $d_{rgb}\in\{16,32,64,128\}$. We observe that a very small color feature dimension (e.g., $d_{rgb}=16$) degrades performance due to limited semantic representational capacity, while excessively large dimensions do not yield significant performance gains. Therefore, to balance inference efficiency with registration precision, we set $d_{rgb}=64$ as our default setting. Additionally, in the fifth block of Table V, we investigate performance variations with different fusion weights $\omega\in\{0.0,0.25,0.50,0.75,1.0\}$, where a larger $\omega$ places more emphasis on the geometric descriptors (see Eq. 12). Our results indicate that both overly high $\omega$ (which overemphasizes geometry) and overly low $\omega$ (which overemphasizes color) lead to degraded registration accuracy. By contrast, a balanced weight (e.g., $\omega=0.50$) achieves higher performance. As a result, we adopt $\omega=0.50$ as our default hyperparameter configuration.
@@
XYZ-RGB Fusion. We finally evaluate the effectiveness of the XYZ-RGB fusion (see Sec. 3.6). We replace the real color point clouds (with actual RGB values) used by ColorPCR [30] with our generated color point clouds (with synthesized RGB values), forming Generative ColorPCR for 3D matching. The third block in Table V demonstrates that, on the 3DMatch dataset, Generative ColorPCR with the synthesized color even outperforms the original ColorPCR with the real color. This advantage is attributed not only to the high-quality pairwise image generation provided by our Match-ControlNet, but also to several key benefits of our generated XYZ-RGB data over real XYZ-RGB data: (i) Mitigating calibration errors: As shown in Fig. 9 (left), some real RGB-D data would suffer from calibration errors, which may lead to misalignment in the colored point clouds. By contrast, our framework, benefiting from the powerful 2D-3D consistency generation ability, effectively reduces such calibration errors, producing higher-quality colored point clouds and enabling more accurate matching. (ii) Mitigating lighting challenges: Fig. 9 (right) shows that RGB images from real-world conditions can degrade under poor lighting, negatively impacting RGB-D point cloud matching performance. By contrast, our generative point cloud registration framework potentially produces images with consistent lighting conditions, independent of real-world illumination variations, thereby enhancing the overall robustness of our method to lighting challenges.
Figure 8: Qualitative comparions between Predator and Generative Predator on Dur360BEV [52] benchmark dataset.
Figure 9: Our Match-ControlNet effectively mitigates calibration errors and lighting challenges commonly encountered in real-world RGB-D data, thereby improving the matching precision of color point cloud registration methods. For example, Generative ColorPCR using synthetic RGB data even outperforms the original ColorPCR trained with real RGB data, as reported in the third block of Table V.